# Measuring gravitational lensing time delays with quantum information processing

### v1

@@ V.2 Quantum Fourier transform
$\quad\mathopen{}\mathclose{{\left(1+e^{\mathrm{i}\omega\Delta t}e^{-2\pi\mathrm{i}\Delta t(k+f_{\mathrm{alias}}T)/T}}}\right).$
@@
$\cdot\mathopen{}\mathclose{{\left(1+\cos(\Delta t(\omega-2\pi f_{\mathrm{alias}}-2\pi k/T))}}\right).$
@@ V.3 Connection to the dihedral hidden subgroup problem
It is well known that the DHSP reduces to a quantum state learning problem called the dihedral coset problem (DCP). (See Ref. Childs and Van Dam (2010) for a comprehensive introduction to the DHSP and the DCP.) In the DCP of size $N$, we are given multiple samples of a quantum coset state
@@
where $a\in\{0,1,\dots,N-1\}$ is uniformly random for each sample, and $l\in\{0,1,\dots,N-1\}$ is a fixed unknown parameter. The goal of the DCP is to find the value of $l$ using the given states. The sample complexity of the DCP is the number of states required to determine $l$ with bounded error, and the computational complexity is the amount of computation needed to process the states and learn $l$. Both complexities are typically analyzed in terms of their asymptotic scaling with $N$. An early result of Ettinger and Høyer showed that the sample complexity of DCP is $O(\log N)$ Ettinger and Høyer (2000) (indeed, an analysis of the optimal recovery procedure shows the sample complexity is $\Theta(\log N)$ Bacon et al. (2006)). However, no known quantum algorithm for the DCP is efficient (i.e., runs in time $\operatorname{poly}(\log N)$), and indeed, there is no known efficient classical or quantum algorithm for the DHSP. Indeed, the belief that no such algorithm exists underlies the presumed security of lattice-based public-key cryptography Regev (2004).
@@
where we map between quantities in the DCP and those in the delay-finding problem as follows: $\frac{t_{0}}{t_{c}}\mapsto a$, $\frac{\Delta t}{t_{c}}\mapsto l$, $\frac{T}{t_{c}}=n_{s}\mapsto N$. Such states can be obtained by measuring the first qubit of the DCP states in the $X$ basis and post-selecting on +1+1 outcomes. This postselection succeeds with probability $1/2$, resulting in only a factor-of-$2$ overhead in the production of time-delay states from dihedral coset states. Thus an efficient algorithm for solving the delay-finding problem can be used to efficiently solve the DCP. ∎
Our approach to solving the delay-finding problem follows the same strategy as in Ettinger and Høyer’s procedure for solving the DHSP by producing coset states, measuring them in an appropriate basis, and inferring $l$ from the results Ettinger and Høyer (2000). While it uses polynomially many samples, this procedure is computationally inefficient, using exponential (in $\log N$) processing. Fortunately, the corresponding $O(T/t_{c})$ classical data processing cost is acceptable for the delay-finding problem, since $T$ and $t_{c}$ are both constants for a certain observation, while the sample complexity is the real bottleneck in microlensing observation.
@@ VI Experimental realization
We first review the resolution and bandwidth of existing approaches to single-photon spectrometry using direct frequency readout without digitization (Section VI.1). These are the only experimental building block needed for Algorithm 1. For the quantum undersampling algorithm (Algorithm 2), we present two schemes. Both schemes are digital and operate with a classical switch acting on the digitization timescale $t_{c}/10$. First, we present a linear-optics implementation for the discrete (quantum) Fourier transform of the photonic state, albeit in practice some of the components are used in quantum optics (Section VI.2). Second, we point out that the state of the light can be transferred to quantum memories, followed by quantum computation (Section VI.3). Although the experimental realization of high-quality quantum memories is challenging in the near term, the advantages of the approach in Section VI.3 compared to that in Section VI.2 are longer possible storage times of the light (compared to delays achievable with delay lines in Section VI.2) and hence larger allowed values of $T$, and an exponential reduction in the number of gates (compared to the number of beam splitters in Section VI.2) and memory (compared to the number of delay lines in Section VI.2) when binary encoding is available. The approach in Section VI.3 is particularly suited for the application of our algorithm in telescope arrays (Section IX) in order to minimize entanglement consumption.
@@ VI.1 Single-photon spectrometry
As a popular scheme for single-photon spectrometry, frequency-to-time mapping can achieve $10\,\mathrm{GHz}$-level resolution with $10^{12}\,\mathrm{Hz}$-level bandwidth via chirped fiber Bragg gratings Davis et al. (2017) or integrated thin-film lithium niobate phase modulators Zhu et al. (2022). The time lens Mazelanik et al. (2020); Joshi et al. (2022) is also capable of transferring frequency information into temporal information. Remarkably, Ref. Mazelanik et al. (2020) employs a spin-wave modulation method and a gradient echo memory to achieve $20\,\mathrm{kHz}$ resolution with $\mathrm{MHz}$-level bandwidth. Next, on-chip spectrometers based on superconducting nanowire single-photon detectors Cheng et al. (2019); Xiao et al. (2022); Kahl et al. (2017) support broadband input with $10^{14}\,\mathrm{Hz}$-level bandwidth with typical spectral resolution around $100\,\mathrm{GHz}$. Additionally, the dual-comb Coddington et al. (2016); Picqué and Hänsch (2020); Zhong et al. (2025) approach has been shown to be powerful in implementing single-photon spectrometry: Ref. Xu et al. (2024) demonstrates $200\,\mathrm{MHz}$-level resolution with $50\,\mathrm{GHz}$ bandwidth; Ref. Peng et al. (2025) provides $125\,\mathrm{MHz}$ resolution with $\sim 10\,\mathrm{GHz}$ bandwidth. There is also a frequency-to-space mapping scheme Nagoro et al. (2025) that achieves $120\,\mathrm{MHz}$ resolution with $15\,\mathrm{GHz}$ bandwidth using a single-photon avalanche diode array.
@@
Finally, we briefly discuss another possible approach to realize high-resolution spectrometry based on ensemble-based quantum memories Heshami et al. (2016); Shinbrough et al. (2023). First, if one uses quantum memories based on inhomogeneous broadening Afzelius et al. (2009); Ma et al. (2021); Mejia et al. (2025), one might be able to measure the frequency of the incoming light by projectively measuring which broadening class of atoms the light is stored in. Second, if the frequency of light is mapped to spatial frequency of the ensemble memory, one possible realization of frequency readout is putting a cavity around the memory and coupling the desired spatial frequency components to the cavity sequentially one by one. We leave the implementation of these ensemble-memory-based spectrometry approaches, along with improving existing ones, as potential future directions of research.
@@ VI.2 Linear optics
We sketch a possible implementation with linear optics (Fig. 4). Incoming light from the source is collected by a telescope. It is routed onto different paths by a classical switch: e.g., movable mirrors that reflect the light. The switch operates on the digitization timescale $\tau_{s}=t_{c}/10$ set by the coherence time of the light, such that the different paths have $\Theta(1)$ amplitude of interference. Delay lines are introduced for each path such that they arrive at an interferometer at the same time. The delay can be realized by additional fiber that the light needs to traverse, or by an atomic cloud with a high index of refraction. The interference is done by a network of beamsplitters and phase shifters comprising a $P$-port, which can realize the QFT unitary of dimension $P$ Reck et al. (1994). Given the digitization time $O(1/t_{c})$ and the total observation time $T$, the number of ports is $P=O(T/t_{c})$, and the number of beam splitters is $O((T/t_{c})^{2})$. Implementation of the fast Fourier transform reduces the count to $O((T/t_{c})\log(T/t_{c}))$ Barak and Ben-Aryeh (2007); Hillerkuss et al. (2010). Finally, the intensity of the light at the output of the $P$-port is measured. Since the number of incident photons is small compared to the number of ports, single-photon detectors are necessary and sufficient. Furthermore, the large number of elements and the associated precision, along with the limited delay time achievable in practice (about a microsecond Guo et al. (2025)), can make implementation challenging.
@@ VI.3 Quantum computing
As a potentially more efficient approach to implement the discretization process and the quantum Fourier transform of Algorithm 2, storage of light in digital quantum memory followed by digital quantum computation was previously considered in the context of quantum telescope arrays Khabiboulline et al. (2019a). In particular, an incoming photon in a superposition of $O(T/t_{c})$ arrival times is coherently stored in $O(\log(T/t_{c}))$ qubits. An incoming photon ($\gamma$) is mapped to quantum memory ($a$) in the following way:
@@
where $\bar{t}$ denotes the binary representation of $t$. In contrast to collapsing the temporal superposition and performing a Fourier transform on the spatial coherences, as in Ref. Khabiboulline et al. (2019b), now we directly apply a quantum Fourier transform on the quantum memory. Readout of the qubits in the computational basis completes the quantum portion of Algorithm 2.
The state transfer from light to quantum memory may be realized with entangling gates between flying photons and qubits, mediated by cavities Duan and Kimble (2004), and non-destructive photon measurements Khabiboulline et al. (2019b). Despite having exponentially-reduced resource scaling with $T/t_{c}$, the experimental realization of such a binary-encoded quantum memory remains challenging.
@@ VII Observation proposal: M dwarf flares
$t_{E}$ $=\frac{\sqrt{4GMD_{L}(1-\frac{D_{L}}{D_{S}})}}{v_{T}c}$ (89)
$\approx 4\,\text{days}\,\mathopen{}\mathclose{{\left(\frac{M}{M_{\text{Jup}}}}}\right)^{1/2}$
@@ VII.1 Example setup
Despite their small size, M dwarfs are one of the most active stellar types, producing flares that last on the order 1101-10 minutes and release energy in the range $10^{28}-10^{34}$ ergs. Flare temperatures are difficult to constrain without multiwavelength spectra. As such, flare temperatures are usually assumed to be $T_{\text{flare}}=9000$ K despite evidence that typical flare temperatures may actually be closer to 11,000 K Jackman et al. (2023), with some rare superflares even reaching peak temperatures of $\gtrsim 15,000$ K Howard et al. (2020).
For the sake of specificity, we will consider a fiducial M5V dwarf  [^1] The ‘5’ in M5V indicates a spectral subclass, specifying the star’s temperature within the M-type sequence (a lower number indicates a hotter star). The ‘V’ signifies luminosity class five, denoting a main-sequence star that generates energy through hydrogen fusion in its core. in what follows. We set the source parameters as $R_{S}=0.2\,R_{\odot}$, an effective temperature $T=3060$ K, and a flare frequency distribution that falls off with $\tilde{\nu}=(3\,\text{day}^{-1})(\frac{E}{10^{30}\,\text{erg}})^{-0.65}$, where $\tilde{\nu}$ is the cumulative frequency of flares occurring per unit time for flares with energy $\geq E$ Paudel et al. (2024). This power law has a range of validity of $\approx 10^{29}-10^{32}\,\mathrm{erg}$.
The physical processes that give rise to flares on M dwarfs are not fully understood, but are believed to be related to magnetic reconnection events in active regions. Both the size and temperature of these regions are not well known, though energetics arguments and the spectra of observed flares suggest emission regions on the order of $2\times 10^{9}\,\mathrm{cm}$ (or $\approx 3R_{\oplus}$) Paudel et al. (2024) for an assumed flare temperature of 9000 K. However, if the flare is actually at higher temperature, then a smaller region can still reproduce the observed energy. We see that, as a result, smaller, hotter flares are more likely to satisfy the constraint in Eq. (14), hence are better targets for our lensing scenario. However, applying our algorithm to larger, cooler flares that do not satisfy the finite-source constraint still yields interesting results; since such events would not have detectable time delays, a non-detection for a given flare places constraints on the spatial size of the flare emission region.
@@
$T_{\text{flare}}$ 10,000 K Temperature of flare
$\tau_{\text{flare}}$ 1 min Flare duration
@@
We wish to explore the parameter space in which stellar flares may be a viable target for such an experiment. As described above, we adopt a fiducial M5V dwarf as our target, situating it in the Galactic Bulge at $D_{S}=8\,\mathrm{kpc}$. We assume a flare temperature $T_{\text{flare}}$ of $10,000\,\mathrm{K}$, a flare size $R_{\text{flare}}$ of $3000\,\mathrm{km}$, and a flare duration $\tau_{\text{flare}}$ of 1 minute. This corresponds to a flare with total energy $\approx 10^{31}\,\mathrm{erg}$, which, given our fiducial flare frequency distribution, would occur at a rate of $\approx 0.67/$day. To compute the number of signal/background photons incident on our telescope per flare, we integrate over the flare/dwarf spectrum (modeling both as blackbodies at their respective temperatures) in the passband ranging from $\lambda_{\text{max}}=510\,\mathrm{nm}$, the peak wavelength of the flare blackbody spectrum, to $\lambda_{\text{min}}=365\,\mathrm{nm}$, the cutoff wavelength determined by Eq. (14), with an additional factor $\varepsilon=0.2$, a heuristically determined prefactor necessary for finite-source effects to be negligible. We summarize fiducial model parameters in Table 1.
Additionally, we account for the effects of dust and atmospheric extinction in the following way. We calculate the dust extinction using the observationally-derived extinction coefficients towards the Galactic Bulge found by Ref. Saha et al. (2019), interpolating logarithmically between the values displayed in their Fig. 8. We set the color excess between $r$ and $z$-band filters, E(rz)E(r-z), to 0.5, as taken from the results of Ref. Saha et al. (2019) in Baade’s window. We find that the resulting $A_{v}$ is consistent with the findings of Ref. Stanek (1996). With the interpolated $A_{\lambda}(f)$ from their Fig. 8 with E(rz)=0.5E(r-z)=0.5, we convert to optical depth as $\tau_{\text{dust}}(f)=A_{\lambda}(f)/1.086$. Additionally, we include the effects of atmospheric extinction using the fiducial values provided for Mauna Kea in the Gemini Observer’s Guide Gem (2024), which result in a typical suppression of $\approx 20\%$ in our passband. The total optical depth is therefore $\tau(f)=\tau_{\text{dust}}(f)+\tau_{\text{atmo}}(f)$.
The resulting expressions for the number of signal (background) photons per fiducial flare $n_{\text{sig}}$ ($n_{\text{bg}}$) are given by
@@
Our technique performs best when telescope collecting area is large. However, we find that often, due to the limited number of photons received per flare, a single flare is insufficient to measure the lens mass with 95% confidence. As a result, we must combine measurements across multiple flares to make a confident measurement. In order to achieve this, we modify Algorithm 1 and Algorithm 2 to propose a new algorithm, which is elaborated in Section VII.2. Fortunately, the total number of photons needed to measure the mean value of $\Delta t$ to $t_{c}$ precision with $95\%$ confidence is still $O(\log(T/t_{c}))$ when the number of photons per flare is not too small. We defer the technical details of the derivation to Section VII.2.
@@
We present numerical simulation results for our delay-finding algorithm for a single flare in Fig. 5. The vertical axis is the confidence level of the delay measurement achieved for one flare that produces $n_{\text{sig}}$ photons, and the horizontal axis is the number of signal photons received per flare. We see that in our fiducial case, a single flare would need to yield $\approx 200$ photons collected in our telescope to achieve a 50% confidence detection, and $\approx 375$ to achieve a 95% confidence detection. Note that we have also included a zero-background limit ($Q=1$, black), which while not applicable to the flare scenario, is more broadly applicable to possible other isolated sources. We see that if an isolated source satisfied the finite-source condition, it would require collecting only 150 total photons from that source to measure the time delay at 95% confidence. This impressive sensitivity to even very faint sources may have broader applications beyond the flare scenario studied here.
@@
Figure 5: This figure shows, with two possible values of the signal-to-background ratio $Q$, how the confidence level of $\Delta t$ measurements increases with the number of signal photons. The confidence level for each $n_{\mathrm{sig}}$ is computed by numerical simulation of a scenario with $A=1.34$ and $T/t_{c}=10^{4}$. Here $Q=1$ corresponds to the scenario with no background or noise photons, and $Q=0.4$ corresponds to the fiducial example of M-dwarf flares considered in our observation proposal. The $95\%$ confidence and $50\%$ confidence are marked by horizontal dashed lines.
@@
The simulation results of multiple-flare combination are shown in Fig. 6. These plots show how the number of flares necessary to provide a detection changes depending on number of photons per flare $n_{\text{sig}}$, signal-to-background parameter $Q\coloneqq n_{\text{sig}}/(n_{\text{sig}}+n_{\text{bg}})$, confidence level, and ratio of maximum time delay to coherence time, $T/t_{c}$. Black and red curves correspond to the analytic bound derived in Section VII.2, which is sufficient but not necessarily optimal. Black and red points correspond to the results of numerical simulations, which provide a much better estimate of the true number of flares needed.

### v2

@@ V.2 Quantum Fourier transform
$\quad\mathopen{}\mathclose{{\left(1+e^{\mathrm{i}\omega_{0}\Delta t}e^{-2\pi\mathrm{i}\Delta t(k+f_{\mathrm{alias}}T)/T}}}\right).$
@@
$\cdot\mathopen{}\mathclose{{\left(1+\cos(\Delta t(\omega_{0}-2\pi f_{\mathrm{alias}}-2\pi k/T))}}\right).$
@@ V.3 Connection to the dihedral hidden subgroup problem
It is well known that the DHSP reduces to a quantum state learning problem called the dihedral coset problem (DCP). (See Ref. [63] for a comprehensive introduction to the DHSP and the DCP.) In the DCP of size $N$, we are given multiple samples of a quantum coset state
@@
where $a\in\{0,1,\dots,N-1\}$ is uniformly random for each sample, and $l\in\{0,1,\dots,N-1\}$ is a fixed unknown parameter. The goal of the DCP is to find the value of $l$ using the given states. The sample complexity of the DCP is the number of states required to determine $l$ with bounded error, and the computational complexity is the amount of computation needed to process the states and learn $l$. Both complexities are typically analyzed in terms of their asymptotic scaling with $N$. An early result of Ettinger and Høyer showed that the sample complexity of DCP is $O(\log N)$ [33] (indeed, an analysis of the optimal recovery procedure shows the sample complexity is $\Theta(\log N)$ [64]). However, no known quantum algorithm for the DCP is efficient (i.e., runs in time $\operatorname{poly}(\log N)$), and indeed, there is no known efficient classical or quantum algorithm for the DHSP. Indeed, the belief that no such algorithm exists underlies the presumed security of lattice-based public-key cryptography [65].
@@
where we map between quantities in the DCP and those in the delay-finding problem as follows: $\frac{t_{0}}{t^{\prime}_{c}}\mapsto a$, $\frac{\Delta t}{t^{\prime}_{c}}\mapsto l$, $\frac{T}{t^{\prime}_{c}}=n_{s}\mapsto N$. Such states can be obtained by measuring the first qubit of the DCP states in the $X$ basis and post-selecting on +1+1 outcomes. This postselection succeeds with probability $1/2$, resulting in only a factor-of-$2$ overhead in the production of time-delay states from dihedral coset states. Thus an efficient algorithm for solving the delay-finding problem can be used to efficiently solve the DCP. ∎
Our approach to solving the delay-finding problem follows the same strategy as in Ettinger and Høyer’s procedure for solving the DHSP by producing coset states, measuring them in an appropriate basis, and inferring $l$ from the results [33]. While it uses polynomially many samples, this procedure is computationally inefficient, using exponential (in $\log N$) processing. Fortunately, the corresponding $\operatorname{poly}(T/t_{c})$ classical data processing cost is acceptable for the delay-finding problem, since $T$ and $t_{c}$ are both constants for a certain observation, while the sample complexity is the real bottleneck in microlensing observation.
@@ VI Experimental realization
We first review the resolution and bandwidth of existing approaches to single-photon spectrometry using direct frequency readout without digitization (Section VI.1). These are the only experimental building blocks needed for Algorithm 1. For the quantum undersampling algorithm (Algorithm 2), we present two schemes. Both schemes are digital and operate with a classical switch acting on the digitization timescale $t^{\prime}_{c}/10$. First, we present a linear-optics implementation for the discrete (quantum) Fourier transform of the photonic state, albeit in practice some of the components are used in quantum optics (Section VI.2). Second, we point out that the state of the light can be transferred to quantum memories, followed by quantum computation (Section VI.3). Although the experimental realization of high-quality quantum memories is challenging in the near term, the advantages of the approach in Section VI.3 compared to that in Section VI.2 are longer possible storage times of the light (compared to delays achievable with delay lines in Section VI.2) and hence larger allowed values of $T$, and an exponential reduction in the number of gates (compared to the number of beam splitters in Section VI.2) and memory (compared to the number of delay lines in Section VI.2) when binary encoding is available. The approach in Section VI.3 is particularly suited for the application of our algorithm in telescope arrays (Section IX) in order to minimize entanglement consumption.
@@ VI.1 Single-photon spectrometry
As a popular scheme for single-photon spectrometry, frequency-to-time mapping can achieve $10\,\mathrm{GHz}$-level resolution with $10^{12}\,\mathrm{Hz}$-level bandwidth via chirped fiber Bragg gratings [66] or integrated thin-film lithium niobate phase modulators [67]. The time lens [38, 39] is also capable of transferring frequency information into temporal information. Remarkably, Ref. [38] employs a spin-wave modulation method and a gradient echo memory to achieve $20\,\mathrm{kHz}$ resolution with $\mathrm{MHz}$-level bandwidth. Next, on-chip spectrometers based on superconducting nanowire single-photon detectors [68, 69, 70] support broadband input with $10^{14}\,\mathrm{Hz}$-level bandwidth with typical spectral resolution around $100\,\mathrm{GHz}$. Additionally, the dual-comb [34, 35, 71] approach has been shown to be powerful in implementing single-photon spectrometry: Ref. [36] demonstrates $200\,\mathrm{MHz}$-level resolution with $50\,\mathrm{GHz}$ bandwidth; Ref. [37] provides $125\,\mathrm{MHz}$ resolution with $\sim 10\,\mathrm{GHz}$ bandwidth. There is also a frequency-to-space mapping scheme [72] that achieves $120\,\mathrm{MHz}$ resolution with $15\,\mathrm{GHz}$ bandwidth using a single-photon avalanche diode array.
@@
Finally, we briefly discuss another possible approach to realize high-resolution spectrometry based on ensemble-based quantum memories [73, 74]. First, if one uses quantum memories based on inhomogeneous broadening [75, 76, 77], one might be able to measure the frequency of the incoming light by projectively measuring which broadening class of atoms the light is stored in. Second, if the frequency of light is mapped to spatial frequency of the ensemble memory, one possible realization of frequency readout is putting a cavity around the memory and coupling the desired spatial frequency components to the cavity sequentially one by one. We leave the implementation of these ensemble-memory-based spectrometry approaches, along with improving existing ones, as potential future directions of research.
@@ VI.2 Linear optics
We sketch a possible implementation with linear optics (Fig. 4). Incoming light from the source is collected by a telescope. It is routed onto different paths by a classical switch: e.g., movable mirrors that reflect the light. The switch operates on the digitization timescale $\tau_{s}=t^{\prime}_{c}/10$ set by the coherence time of the light, such that the different paths have $\Theta(1)$ amplitude of interference. Delay lines are introduced for each path such that they arrive at an interferometer at the same time. The delay can be realized by additional fiber that the light needs to traverse, or by an atomic cloud with a high index of refraction. The interference is done by a network of beamsplitters and phase shifters comprising a $P$-port, which can realize the QFT unitary of dimension $P$ [78]. Given the digitization time $O(1/t^{\prime}_{c})$ and the total observation time $T$, the number of ports is $P=O(T/t^{\prime}_{c})$, and the number of beam splitters is $O((T/t^{\prime}_{c})^{2})$. Implementation of the fast Fourier transform reduces the count to $O((T/t^{\prime}_{c})\log(T/t^{\prime}_{c}))$ [79, 80]. Finally, the intensity of the light at the output of the $P$-port is measured. Since the number of incident photons is small compared to the number of ports, single-photon detectors are necessary and sufficient. Furthermore, the large number of elements and the associated precision, along with the limited delay time achievable in practice (about a microsecond [81]), can make implementation challenging.
@@ VI.3 Quantum computing
As a potentially more efficient approach to implement the discretization process and the quantum Fourier transform of Algorithm 2, storage of light in digital quantum memory followed by digital quantum computation was previously considered in the context of quantum telescope arrays [5]. In particular, an incoming photon in a superposition of $O(T/t^{\prime}_{c})$ arrival times is coherently stored in $O(\log(T/t^{\prime}_{c}))$ qubits. An incoming photon ($\gamma$) is mapped to quantum memory ($a$) in the following way:
@@
where $\bar{t}$ denotes the binary representation of $t$. In contrast to collapsing the temporal superposition and performing a Fourier transform on the spatial coherences, as in Ref. [6], now we directly apply a quantum Fourier transform on the quantum memory. Readout of the qubits in the computational basis completes the quantum portion of Algorithm 2.
The state transfer from light to quantum memory may be realized with entangling gates between flying photons and qubits, mediated by cavities [82], and non-destructive photon measurements [6]. Despite having exponentially-reduced resource scaling with $T/t^{\prime}_{c}$, the experimental realization of such a binary-encoded quantum memory remains challenging.
@@ VII Observation proposal: M dwarf flares
$t_{E}$ $=\frac{\sqrt{4GMD_{L}(1-\frac{D_{L}}{D_{S}})}}{v_{T}c}\approx 4\,\text{days}\,\mathopen{}\mathclose{{\left(\frac{M}{M_{\text{Jup}}}}}\right)^{1/2}$ (92)
@@ VII.1 Example setup
Despite their small size, M dwarfs are one of the most active stellar types, producing flares that last on the order 1101-10 minutes and release energy in the range $10^{28}\,\mathrm{ergs}-10^{34}\,\mathrm{ergs}$. Flare temperatures are difficult to constrain without multiwavelength spectra. As such, flare temperatures are usually assumed to be $T_{\text{flare}}=9\,000$ K despite evidence that typical flare temperatures may actually be closer to 11 $1\,000\,\mathrm{K}$ [83], with some rare superflares even reaching peak temperatures of $\gtrsim 15\,000$ K [84].
For the sake of specificity, we will consider a fiducial M5V dwarf  [^2] The ‘5’ in M5V indicates a spectral subclass, specifying the star’s temperature within the M-type sequence (a lower number indicates a hotter star). The ‘V’ signifies luminosity class five, denoting a main-sequence star that generates energy through hydrogen fusion in its core. in what follows. We set the source parameters as $R_{S}=0.2\,R_{\odot}$, an effective temperature $T=3060$ K, and a flare frequency distribution that falls off with $\tilde{\nu}=(3\,\text{day}^{-1})(\frac{E}{10^{30}\,\text{ergs}})^{-0.65}$, where $\tilde{\nu}$ is the cumulative frequency of flares occurring per unit time for flares with energy $\geq E$ [86]. This power law has a range of validity of $\approx 10^{29}\,\mathrm{erg}-10^{32}\,\mathrm{erg}$.
The physical processes that give rise to flares on M dwarfs are not fully understood, but are believed to be related to magnetic reconnection events in active regions. Both the size and temperature of these regions are not well known, though energetics arguments and the spectra of observed flares suggest emission regions on the order of $2\times 10^{9}\,\mathrm{cm}$ (or $\approx 3R_{\oplus}$) [86] for an assumed flare temperature of 9000 K. However, if the flare is actually at higher temperature, then a smaller region can still reproduce the observed energy. We see that, as a result, smaller, hotter flares are more likely to satisfy the constraint in Eq. (14), hence are better targets for our lensing scenario. However, applying our algorithm to larger, cooler flares that do not satisfy the finite-source constraint still yields interesting results; since such events would not have detectable time delays, a non-detection for a given flare places constraints on the spatial size of the flare emission region.
@@
$T_{\text{flare}}$ $10\,000\,\mathrm{K}$ Temperature of flare
$\tau_{\text{flare}}$ 2 min Flare duration
@@
We wish to explore the parameter space in which stellar flares may be a viable target for such an experiment. As described above, we adopt a fiducial M5V dwarf as our target, situating it in the Galactic Bulge at $D_{S}=8\,\mathrm{kpc}$. We assume a flare temperature $T_{\text{flare}}$ of $10\,000\,\mathrm{K}$, a flare size $R_{\text{flare}}$ of 3000 km, and a flare duration $\tau_{\text{flare}}$ of 1 minute. This corresponds to a flare with total energy $\approx 10^{31}\,\mathrm{erg}$, which, given our fiducial flare frequency distribution, would occur at a rate of $\approx 0.67/$day. To compute the number of signal/background photons incident on our telescope per flare, we integrate over the flare/dwarf spectrum (modeling both as blackbodies at their respective temperatures) in the passband ranging from $\lambda_{\text{max}}=510\,\mathrm{nm}$, the peak wavelength of the flare blackbody spectrum, to $\lambda_{\text{min}}=365\,\mathrm{nm}$, the cutoff wavelength determined by Eq. (14), with an additional factor $\varepsilon=0.2$, a heuristically determined prefactor necessary for finite-source effects to be negligible. We summarize fiducial model parameters in Table 1.
Additionally, we account for the effects of dust and atmospheric extinction in the following way. We calculate the dust extinction using the observationally-derived extinction coefficients towards the Galactic Bulge found by Ref. [87], interpolating logarithmically between the values displayed in their Fig. 8. We set the color excess between $r$ and $z$-band filters, E(rz)E(r-z), to 0.5, as taken from the results of Ref. [87] in Baade’s window. We find that the resulting $A_{v}$ is consistent with the findings of Ref. [88]. With the interpolated $A_{\lambda}(f)$ from their Fig. 8 with E(rz)=0.5E(r-z)=0.5, we convert to optical depth as $\tau_{\text{dust}}(f)=A_{\lambda}(f)/1.086$. Additionally, we include the effects of atmospheric extinction using the fiducial values provided for Mauna Kea in the Gemini Observer’s Guide [89], which result in a typical suppression of $\approx 20\,\%$ in our passband. The total optical depth is therefore $\tau(f)=\tau_{\text{dust}}(f)+\tau_{\text{atmo}}(f)$.
The resulting expressions for the number of signal (background) photons per fiducial flare per minute $n_{\text{sig}}$ ($n_{\text{bg}}$) are given by
@@
Our technique performs best when telescope collecting area is large. However, we find that often, due to the limited number of photons received per flare, a single flare is insufficient to measure the lens mass with $95\,\%$ confidence. As a result, we must combine measurements across multiple flares to make a confident measurement. In order to achieve this, we modify Algorithm 1 and Algorithm 2 to propose a new algorithm, which is elaborated in Section VII.2. Fortunately, the total number of photons needed to measure the mean value of $\Delta t$ to $t_{c}$ precision with $95\,\%$ confidence is still $O(\log(T/t_{c}))$ when the number of photons per flare is not too small. We defer the technical details of the derivation to Section VII.2.
Note that numerical evaluation of $n_{\mathrm{sig}}$ yields up to several hundred signal photons for a typical $1$-minute flare for the largest ground-based telescope in the foreseeable future. Since this is the photon number corresponding to the entire passband from $\lambda_{\mathrm{max}}=510\,\mathrm{nm}$ to $\lambda_{\mathrm{min}}=365\,\mathrm{nm}$, it is a broadband signal with coherence time $t_{c}\sim 10^{-15}\,\mathrm{s}$. According to Section III.7, when $\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}$ are fixed, the size of the search space, $T/t_{c}$, is solely determined by $T$, the upper bound of the time delay.
@@
We present numerical simulation results for our delay-finding algorithm for a single flare in Fig. 5. The vertical axis is the confidence level of the delay measurement achieved for one flare that produces $n_{\text{sig}}$ photons, and the horizontal axis is the number of signal photons received per flare. We see that in our fiducial case, when $T=10^{-9}\,\mathrm{s}$ (corresponding to $T/t_{c}=10^{6}$), a single flare would need to yield $\approx 300$ signal photons collected in our telescope to achieve a $70\,\%$ confidence detection, and $\approx 500$ to achieve a $95\,\%$ confidence detection. For a longer possible time delay, $T=10^{-7}\,\mathrm{s}$ (corresponding to $T/t_{c}=10^{8}$), a single flare with $400$ signal photons can achieve $\geq 60\,\%$ confidence detection. Note that we have also included a zero-background limit ($Q=1$, blue), which while not applicable to the flare scenario, is more broadly applicable to possible other isolated sources. We see that if an isolated source satisfies the finite-source condition, it would require collecting only up to 200 total photons from that source to measure the time delay at $90\,\%$ confidence. This impressive sensitivity to even very faint sources may have broader applications beyond the flare scenario studied here.
@@
Figure 5: This figure shows, with two possible values of the signal-to-background ratio $Q$, how the confidence level of $\Delta t$ measurements increases with the number of signal photons. The confidence level for each $n_{\mathrm{sig}}$ is computed by numerical simulation of a scenario with $A=1.34$. Here $Q=1$ corresponds to the scenario with no background or noise photons, $Q=0.4$ corresponds to the fiducial example of M-dwarf flares considered in our observation proposal, $T/t_{c}=10^{6}$ corresponds to $\Delta t\leq T=10^{-9}\,\mathrm{s}$, and $T/t_{c}=10^{8}$ corresponds to $T=10^{-7}\,\mathrm{s}$. The $90\,\%$ confidence, $70\,\%$ confidence, and $50\,\%$ confidence are marked by horizontal dashed lines.
$T/t_{c}$ $10^{6}$ $10^{7}$ $10^{8}$ $10^{9}$
$M/M_{\mathrm{Jup}}$ $\sim 0.1$ $\sim 1$ $\sim 10$ $\sim 100$
$t_{E}$ (days) $\sim 1.3$ $\sim 4$ $\sim 13$ $\sim 40$
$m_{\mathrm{typ}}$ $\lesssim 1$ $\lesssim 3$ $\lesssim 9$ $\lesssim 27$
$n_{\mathrm{sig}}$ 150 200 400 150 200 400 150 200 400 150 200 400
[…]
Table 2: Summary of the number of flares needed to achieve certain confidence levels. We use the baseline setting discussed in our M-dwarf flare observation proposal: $Q=0.4,A=1.34,t_{c}=10^{-15}\,\mathrm{s}$. For each $T/t_{c}$, we compute its corresponding lens mass $M$ and roughly estimate the event duration $t_{E}$ using Eq. (9) and Eq. (92), respectively. With the assumption that a flare happens every $\sim 1.5$ days on average, we estimate the typical number of flares one can expect for each duration of event, denoted by $m_{\mathrm{typ}}$.
@@
Simulation results for the multiple-flare combination are shown in Table 2. This table shows how the number of flares needed to provide a detection changes with the number of photons per flare $n_{\text{sig}}$, confidence level, and the ratio of maximum time delay to coherence time, $T/t_{c}$.